# Interpretable representation learning of quantum data enabled by probabilistic variational autoencoders

### v2

@@ Appendix E Study of the latent neuron variances for spin models benchmarks
From a practical perspective, the spectral entropy compresses the normalized power spectral density of a signal into a single scalar value, and is computed using the Shannon entropy of its frequency distribution. In this framework, lower spectral entropy corresponds to more structured, periodic, input configurations, while higher values indicate increased stochasticity and disorder. This metric has been employed across various domains, including molecular component identification [49], the characterization of chaotic dynamics [50] and classification tasks [51], demonstrating its broad applicability for quantifying structure and randomness in complex data.

### v3

@@ Appendix F Study of the latent neuron variances for spin models benchmarks
From a practical perspective, the spectral entropy compresses the normalized power spectral density of a signal into a single scalar value, and is computed using the Shannon entropy of its frequency distribution. In this framework, lower spectral entropy corresponds to more structured, periodic, input configurations, while higher values indicate increased stochasticity and disorder. This metric has been employed across various domains, including molecular component identification [li2021spectral], the characterization of chaotic dynamics [xiong2021spectral] and classification tasks [swetapadma2023novel], demonstrating its broad applicability for quantifying structure and randomness in complex data.